A television commercial advertises that a certain type of light bulb will last on average, 200 hours longer that three times the life of another type of light bulb. If consumers tests show that the advertised bulb lasts 4700 hours, how many hours must the other type of bulb last for the advertisers claim to be valid?
Answer by Theo(13342)

the advertised bulb lasts 4700 hours.

this is 200 hours longer than three times the life of another type bulb.

let x equal the life of the other type bulb.

your equation becomes 4700 = 3 * x + 200.

subtract 200 from both sides of this equation to get 4500 = 3 * x.

divide both sides of this equation by 3 to get 1500 = x.

the life of the other type bulb is 1500 hours on average.

the advertised bulb last 200 hours more than 3 times 1500 on average.

that's 200 + 4500 = 4700 hours on average.

this confirms the solution is correct.

the solution is that the other type bulb must last 1500 hours on average for the advertisers claim to be valid.